Question 5

Demand in a market dominated by two firms (a Cournot duopoly) is determined according to: P = 200 – 2(Q1 + Q2), where P is the market price, Q1 is the quantity demanded by Firm 1, and Q2 is the quantity demanded by Firm 2. The marginal cost and average cost for each firm is constant; AC=MC = $60.

The cournot-duopoly equilibrium profit for each firm is _____.

Hint: Write your answer to two decimal places.

QUESTION 6

  1. The following matrix shows the pricing strategies and resultant profits for two profit-maximizing firms. The profit's are written  (Z,Y) where Z is the profit that Firm A (the firm that chooses which row) receives and Y is the profit Firm B (the firm that chooses the column) receives.

    Strategies High Low
    High X,X 4,11.5
    Low 11.5,4 10,10

    What is the lowest value of X such that High, High is a Nash Equilibrium in this scenario.

    Hint: Write your answer to two decimal places.

    Hint two: X=35 would make High, High a Nash Equilibrium since neither firm would benefit from deviating. However, I want to know the smallest value of X such that High, High is the Nash Equilibrium.

10 points   

QUESTION 7

  1. Demand in a market dominated by two firms (a Cournot duopoly) is determined according to: P = 200 – 2(Q1 + Q2), where P is the market price, Q1 is the quantity demanded by Firm 1, and Q2 is the quantity demanded by Firm 2. The marginal cost and average cost for each firm is constant; AC=MC = $64.

    The cournot-duopoly equilibrium quantity produced by each firm is _____.

    Hint: Write your answer to two decimal places.
Answer #1

Question 5:
Each firm maximizes profit according to the rule: MR = MC.
P = 200 - 2(Q1 + Q2) = 200 - 2Q1 - 2Q2
Firm 1: Total Revenue, TR1 = P*Q1 = (200 - 2Q1 - 2Q2)*Q1 = 200Q1 - 2Q12 - 2Q1Q2
Marginal Revenue, MR1 = TR 200 200-4Q1-2072
So, MR1 = MC gives,
200 - 4Q1 - 2Q2 = 60
So, 4Q1 = 200 - 60 - 2Q2 = 140 - 2Q2
So, Q1 = (140/4) - (2Q2/4)
So, Q1 = 35 - 0.5Q2
This is the best response function of firm 1, BR1.
As demand function and MC are same for both firms, so best response function of firm 2 is:
Q2 = 35 - 0.5Q1
Now, substituting BR1 in BR2, we get,
Q2 = 35 - 0.5(35 - 0.5Q2) = 35 - 17.5 + 0.25Q2
So, Q2 - 0.25Q2 = 17.5
So, 0.75Q2 = 17.5
So, Q2 = 17.5/0.75 = 23.33
So, Q1 = 35 - 0.5Q2 = 35 - 0.5(23.33) = 35 - 11.67 = 23.33
So, P = 200 – 2(Q1 + Q2) = 200 – 2(23.33 + 23.33) = 200 - 93.32 = 106.68

So, equilibrium profit for each firm = Total revenue for each firm - Total cost of each firm = P*Q1 - AC*Q1 = (P - AC)*Q1
= (106.68 - 60)*23.33 = $1,089.04

(Note: Post one question at a time.)